# In-context Inverse Optimality for Fair Digital Twins: A Preference-based approach

### v1

@@ Abstract
Digital Twins (DTs) are increasingly used as autonomous decision-makers in complex socio-technical systems. Their mathematically optimal decisions often diverge from human expectations, exposing a persistent gap between algorithmic and bounded human rationality. This work addresses this gap by proposing a framework that operationalizes fairness as a learnable objective within optimization-based Digital Twins. We introduce a preference-driven learning pipeline that infers latent fairness objectives directly from human pairwise preferences over feasible decisions. A novel Siamese neural network is developed to generate convex quadratic cost functions conditioned on contextual information. The resulting surrogate objectives align optimization outcomes with human-perceived fairness while maintaining computational efficiency. The approach is demonstrated on a COVID-19 hospital resource allocation scenario. This study provides an actionable path toward embedding human-centered fairness in the design of autonomous decision-making systems.
@@
Engineering ethics in control and AI; Human-centric automation/AI Systems, and human agency; Cyber-physical social systems (CPSS); Econometric models and methods; Cyber-physical and human systems (CPHS); System dynamics and control in CPHS.
@@ 1 Introduction
Digital Twins (DTs) are increasingly deployed as autonomous decision-makers in complex socio-technical systems, ranging from energy management (Djebali et al., 2024) and mobility planning (Somma et al., 2025) to healthcare logistics (Katsoulakis et al., 2024). In these settings, the digital counterpart continuously observes, predicts, and optimizes the behavior of the physical system, selecting control or planning actions according to a formally defined objective. Yet, while such decision engines are often optimal in a mathematical sense, their recommendations frequently diverge from human expectations. Although they may minimize costs, emissions, or delays, their outputs often remain inequitable, opaque, or simply “unnatural” (Stary, 2021; Schink et al., 2026).
@@
A common remedy has been to design handcrafted utility functions intended to approximate human values. However, such formulations are extremely cumbersome to construct and calibrate. As a concrete example, (Pinto et al., 2025) proposes an optimization-based framework to determine the placement of nature-based interventions (e.g., parks, trees, and green walls). In their formulation, the mixed-integer program’s objective coefficients must encode social, environmental, and aesthetic acceptability. Since these valuations are not analytically defined, the authors hand-tuned convolutional-like kernel weights, trading environmental indicators against perceived social impact. This procedure is error-prone, non-repeatable, and highly domain-sensitive: two solutions that are equally “optimal” in the MILP sense may be perceived very differently by humans, yet nothing in the formulation can express such preference gaps.
@@
Recent advances in preference-based optimization (Previtali et al., 2023) and reinforcement learning from human feedback (Christiano et al., 2017) offer a promising alternative to capture human expectation into decision engines. These approaches have been effectively used to align large language models to human preferences (Rafailov et al., 2024), improve autonomous driving systems (Nagahama et al., 2020), and tune control policies in cyber-physical systems (Zhu et al., 2023). Collectively, these ideas signal a paradigm shift: objectives are no longer designed, but elicited. What remains largely unexplored, however, is how such data-driven preference models can be reintegrated into optimization frameworks to explicitly capture ethical considerations (Quaresmini et al., 2025; Villa et al., 2025b).
@@
Following this trend, rather than viewing fairness as an externally imposed constraint, we interpret it as an emergent property of bounded rational reasoning that needs to be captured by the system without the needs for intermediaries and, therefore, potentially introducing their bias. This reframing has several implications. First, a fair DT must learn what “good” means in a specific context. Second, it highlights the need for procedural fairness–decision processes that involve humans in the loop, allowing their feedback to shape the digital model over time. Finally, the resulting decision problems must remain computationally efficient; user preferences must therefore be modeled through representations that are tractable and compatible with modern numerical optimization solvers (Agrawal et al., 2021; Abdufattokhov et al., 2021).
@@
Building upon these insights, we propose a data-driven methodology that learns fairness as a latent objective inferred from pairwise user preferences. Starting from a classical optimization model, the DT explores feasible decisions and queries human evaluators (or simulated surrogates) to indicate which options they find more acceptable. By maintaining convexity, the learned cost can be seamlessly reinserted into the original solution pipeline with little-to-no additional computational cost, yielding decisions that are both optimal and socially aligned.
1.1 Our Contribution
@@
This work advances fairness-aware Digital Twin design from theoretical framing to practical implementation. Specifically, we contribute a complete end-to-end pipeline that learns and embeds human-aligned objectives within classical optimization models:
@@
    Preference-driven learning pipeline. We develop a data-driven methodology that infers latent fairness objectives directly from human pairwise preferences over feasible solutions.
@@
The remainder of this paper is organized as follows. Section 2 formalizes the proposed preference-based framework for learning surrogate objectives from human feedback. Section 3 is dedicated to present our novel neural architecture. Section 4 presents a proof-of-concept Digital Twin demonstrator for hospital resource allocation, illustrating the method’s practical implementation. Finally, Section 5 summarizes the findings and outlines directions for future research on fairness-aware Digital Twins.
@@ 2 Learning a Surrogate Objective from Preferences
When designing a Digital Twin–regardless of its domain–there is always a decision engine that selects the best action to perform on the actual system according to some user-defined criteria. This engine typically solves an optimization problem of the form:
@@
Although this formulation is mathematically well-posed, it assumes that the objective $f$ is explicitly known. In many practical scenarios, especially those involving fairness, comfort, or social acceptability, such an explicit form is unavailable.
@@ 2.1 Preference-Based Cost Learning
To bridge this gap, we propose a data-driven reformulation in which the cost function itself is learned from human feedback. Instead of explicitly prescribing what the system should optimize, we elicit user preferences between feasible alternatives and infer a latent objective function that reproduces those preferences.
To do so, starting from an instance of (1), we first sample a set of feasible decisions:
@@
In many applications, the user can also be informed of the nominal objective values $f(u_{A},x_{0})$ and $f(u_{B},x_{0})$, which often represent measurable performance indicators such as cost, delay, or negative utility. In the following, we assume that we are in one of such settings.
@@
$\mathcal{D}=\{(x_{0,i},u_{A,i},u_{B,i},\ell_{i})\}_{i=1}^{M}$ (3)
records the empirical preference label $\ell_{i}\in\{-1,+1\}$, where $\ell_{i}=+1$ indicates that $u_{A,i}$ is preferred to $u_{B,i}$.
Given this data, we seek a cost function $f_{\theta}(u,x_{0})$ that pairwise orderings reproduce the observed preferences. In particular, in the remainder of the paper, to do so, we will resort to neural networks. Formally, this learning problem can be posed as the Rank-SVM-like program
$\min_{\theta,\{\varepsilon_{i}\}}$ $\sum_{i}\varepsilon_{i}+\mathcal{R}(f_{\theta})$ (4a)
s.t. $\ell_{i}\!\big(f_{\theta}(u_{A,i},x_{0,i})-f_{\theta}(u_{B,i},x_{0,i})\big)\leq-\delta+\varepsilon_{i},\quad\forall i,$ (4b)
$\varepsilon_{i}\geq 0~~~\forall i,$ (4c)
@@
Once identified, the learned surrogate $f_{\theta}$ replaces the nominal objective in (1), yielding a preference-aware optimization model:
@@ 3 Neural Realization and Training
While (4) defines the desired identification problem, solving it directly would require handling a large number of inequality constraints. To obtain a scalable implementation, we relax the constrained formulation into its differentiable empirical-risk counterpart by replacing each inequality with a hinge ranking loss
$\mathcal{L}_{\mathrm{rank}}=\max\bigl(0,\,f_{\theta}(u_{A},x_{0})-f_{\theta}(u_{B},x_{0})+\delta\bigr),$ (6)
and minimize its empirical expectation over the dataset via stochastic gradient descent. This loss softly enforces the inequality $f_{\theta}(u_{A},x_{0})<f_{\theta}(u_{B},x_{0})-\delta$ whenever $A\succ B$.
To guarantee computational tractability of the downstream optimization, the model $f_{\theta}$ is not learned as a generic neural network but is constrained to produce convex quadratic objectives of the form
@@
where $H(x_{0})\succeq 0$ Abdufattokhov et al. (2021). This choice guarantees that the optimization problem (5) can be solved with standard quadratic ( mixed-integer) solvers without altering the constraint structure or the solver stack and possibly preserving convexity.
To learn the mapping $x_{0}\mapsto(H(x_{0}),q(x_{0}))$, we employ a Siamese training scheme. Two identical copies of the same network $f_{\theta}(\cdot,x_{0})$ are evaluated on the paired actions $(u_{A},x_{0})$ and $(u_{B},x_{0})$, and trained through the hinge ranking loss (6). This architecture is not a different model class but a numerical device that computes both sides of each preference pair through the same parameterized function, ensuring that learning directly enforces the desired pairwise ordering.
@@
Although input–convex neural architectures (ICNNs) could in principle learn more general convex surrogates, optimizing over these functions would render (5) fully non-linear. Although when optimizing over continuous variables, this would not be much of a problem, it might become extremely problematic in mixed integer settings, as MINLP solvers are rare. Our parametrization deliberately avoids this.
@@ 3.1 Reducing Training Complexity via PCA Compression
The overall workflow of the demonstrator is summarized in Fig. 1. It comprises five main stages: (i) generation of diverse feasible decisions via biased optimization; (ii) scoring of these solutions through a simulated user preference model; (iii) construction of a pairwise preference dataset; (iv) learning of a surrogate convex quadratic objective; and (v) reintegration of the learned objective into the original solver for fairness-aware optimization.
Feasible Solution GenerationUser Preference AcquisitionPairwise Dataset ConstructionSiamese DNN Quadratic Cost GeneratorLearned Convex Quadratic Objective $\tfrac{1}{2}u^{\top}H(x_{0})u+q(x_{0})^{\top}u$Reintegration into Solver
Figure 1: Schematic overview of the demonstrator pipeline. Feasible suboptimal solutions are generated via biased optimization, scored by a simulated user function, and used to train a Siamese neural network that learns a convex quadratic objective consistent with user preferences.
@@ 4.6 Experimental Evaluation
5 Conclusions
This paper proposed a data-driven framework to embed fairness in optimization-based Digital Twins through preference learning. We introduced a pipeline that infers human-aligned surrogate objectives from pairwise preferences and realized it via a Siamese neural network that generates convex quadratic costs, ensuring tractability and solver compatibility. A proof-of-concept study on COVID-19 hospital resource allocation demonstrated that the learned objectives can improve perceived fairness without compromising nominal efficiency. Future work will extend this framework to scenarios involving multiple users with potentially diverging preferences, aiming to reconcile conflicting feedback signals and model collective fairness equilibria. Additional research will explore real human-in-the-loop and online learning scenarios.

### v2

# Inverse Optimality for Fair Digital Twins: A Preference-based approach

@@ Abstract
Digital Twins (DTs) are increasingly used as autonomous decision-makers in complex socio-technical systems. However, their mathematically optimal decisions often diverge from human expectations, revealing a persistent mismatch between algorithmic and bounded human rationality. This work addresses this challenge by proposing a framework that introduces fairness as a learnable objective within optimization-based Digital Twins. In this respect, a preference-driven learning workflow that infers latent fairness objectives directly from human pairwise preferences over feasible decisions is introduced. A dedicated Siamese neural network is developed to generate convex quadratic cost functions conditioned on contextual information. The resulting surrogate objectives drive the optimization procedure toward solutions that better reflect human-perceived fairness while maintaining computational efficiency. The effectiveness of the approach is demonstrated on a COVID-19 hospital resource allocation scenario. Overall, this work offers a practical solution to integrate human-centered fairness into the design of autonomous decision-making systems.
@@
Human-centered production and logistics; Human-technology integration in manufacturing; Data-driven and AI-based modelling of production and logistics.
@@ 1 Introduction
Digital Twins (DTs) are increasingly deployed as autonomous decision-makers in complex socio-technical systems, ranging from energy management (Djebali et al., 2024) and mobility planning (Somma et al., 2025) to healthcare logistics (Katsoulakis et al., 2024). In these settings, the digital counterpart continuously observes, predicts and optimizes the behavior of the physical system, selecting control or planning actions according to a formally defined objective. However, while such decisions are often optimal in a mathematical sense, their recommendations frequently diverge from human expectations. Although they may minimize costs, emissions or delays, their outputs often remain inequitable, opaque, or simply “unnatural” to the human thought  (Stary, 2021; Schink et al., 2026).
@@
A common remedy has been to design handcrafted utility functions intended to approximate human values. However, such formulations are extremely cumbersome to construct and calibrate. As a concrete example, (Pinto et al., 2025) proposes an optimization-based framework to determine the placement of nature-based interventions (e.g., parks, trees, and green walls). In their formulation, the mixed-integer program’s objective coefficients must encode social, environmental, and aesthetic acceptability. Since these valuations are not analytically defined, the authors hand-tuned convolutional-like kernel weights, trading environmental indicators against perceived social impact. Clearly, this procedure is error-prone, non-repeatable, and highly domain-sensitive: two solutions that are equally “optimal” in the MILP sense may be perceived very differently by humans, and nothing in the formulation can express such preference gaps.
@@
Other approaches make use of preference-based optimization (Previtali et al., 2023) and reinforcement learning from human feedback (Christiano et al., 2017), and offer a promising alternative to capture human expectation into decision engines. These approaches have been effectively used to align large language models to human preferences (Rafailov et al., 2024), improve autonomous driving systems (Nagahama et al., 2020), and tune control policies in cyber-physical systems (Zhu et al., 2023). Collectively, these ideas signal a paradigm shift: objectives are no longer designed, but elicited. However, how such data-driven preference models can be reintegrated into optimization frameworks to explicitly capture ethical considerations remains largely unexplored (Quaresmini et al., 2025; Villa et al., 2025b).
@@
In this context, rather than viewing fairness (that is the latent context-dependent preference structure that humans implicitly apply when judging alternative decisions, i.e., the set of decision trade-offs that people consistently perceive as equitable or acceptable in a given situation) as an externally imposed constraint, we interpret it as an emergent property of bounded rational reasoning that the system must learn directly, without intermediaries who might potentially introduce bias. This reframing has several implications. First, a fair DT must learn what “good” means in a specific context. Second, it emphasizes the importance of procedural fairness, where humans remain in the loop and iteratively guide the digital model through their feedback. Finally, the resulting decision problems must remain computationally efficient; user preferences must therefore be modeled through representations that are tractable and compatible with modern numerical optimization solvers (Agrawal et al., 2021; Abdufattokhov et al., 2021).
@@
Building upon these insights, we propose a data-driven methodology that learns fairness as a latent objective inferred from pairwise user preferences. Starting from a classical optimization model, the DT explores feasible decisions and queries human evaluators (or simulated surrogates) to indicate which options they find more acceptable. By maintaining convexity, the learned cost can be seamlessly reinserted into the original solution workflow with little-to-no additional computational cost, yielding decisions that are both optimal and socially aligned.
@@
Contribution. This work advances fairness-aware Digital Twin design from theoretical framing to practical implementation. Specifically, we provide a complete end-to-end workflow that learns and embeds human-aligned objectives within classical optimization models. More precisely, we address what follows.
@@
    Preference-driven learning workflow. We develop a data-driven methodology that infers latent fairness objectives directly from human pairwise preferences over feasible solutions.
@@
Paper organization. The remainder of this paper is organized as follows. Section 2 formalizes the proposed preference-based framework for learning surrogate objectives from human feedback. Section 3 is dedicated to present our novel neural architecture. Section 4 presents a proof-of-concept Digital Twin demonstrator for hospital resource allocation, illustrating the method’s practical implementation. Finally, Section 5 summarizes the findings and outlines directions for future research on fairness-aware Digital Twins.
@@ 2 Learning a Surrogate Objective from Preferences
Regardless of its domain, designing a Digital Twin always requires a decision engine that selects the best action to perform on the actual system according to some user-defined criteria. This engine typically solves an optimization problem of the form:
@@
Although this formulation is mathematically well-posed, it assumes that the objective $f$ is explicitly known. In many practical scenarios, especially those involving fairness, comfort, or social acceptability, such an explicit form is however unavailable or very difficult to be obtained, an operation that relies almost entirely on (potentially biased) human operators.
@@ 2.1 Preference-Based Cost Learning
To address this issue, we propose a data-driven reformulation where the cost function itself is learned from human feedback. Instead of explicitly prescribing what the system should optimize, we elicit user preferences between feasible alternatives and infer a latent objective function that, once placed into an appropriate optimization process, would reproduces those preferences. To do so, starting from an instance of (1), we first sample a set of feasible decisions:
@@
In many applications, the user can also be informed of the nominal objective values $f(u_{A},x_{0})$ and $f(u_{B},x_{0})$, which often represent a measurable performance indicator such as cost, delay, or negative utility. In the following, we assume that we fall such setting.
@@
$\mathcal{D}=\{(x_{0,i},u_{A,i},u_{B,i},\ell_{i})\}_{i=1}^{N}$ (3)
records the empirical preference label $\ell_{i}\in\{-1,+1\}$, where $\ell_{i}=+1$ indicates that $u_{A,i}$ is preferred to $u_{B,i}$, that means $f(u_{A_{i}},x_{0})<f(u_{B_{i}},x_{0})$, and vice versa.
Given this data, we seek a cost parametric function $f(u,x_{0})=f_{\theta}(u,x_{0})$, where $\theta=\{\theta_{i}\}$ denotes the set of learnable parameters that characterize the function modeling human preference. To this aim, in the remainder of the paper, we will exploit neural networks. In particular, we consider $f_{\theta}(u,x_{0})$ to be a neural network where $\theta_{i}$ are the weights of the network to be identified. In the next section, we show how $f_{\theta}$ can be learned in a quadratic fashion. Formally, this learning problem can be posed as the Rank-SVM-like program
$\min_{\theta,\{\varepsilon_{i}\}}$ $\sum_{i=1}^{N}\varepsilon_{i}+\mathcal{R}(f_{\theta})$ (4a)
s.t. $\big(f_{\theta}(u_{A,i},x_{0,i})-f_{\theta}(u_{B,i},x_{0,i})\big)\,\ell_{i}\leq-\delta+\varepsilon_{i},$ (4b)
$\varepsilon_{i}\geq 0,\ \forall i,$ (4c)
@@
Once identified, the learned surrogate $f_{\theta}$ is used to parameterize (1) (or replacing the original, biased $f$), yielding a preference-aware optimization model:
@@ 3 Neural Realization and Training
While (4) defines the desired identification problem, solving it directly with standard deep learning toolboxes is difficult. To lessen this issue, we relax the constrained formulation into its differentiable empirical-risk counterpart by replacing each inequality with a hinge ranking loss
$\mathcal{L}_{\mathrm{rank}}=\sum_{i=1}^{N}\max\bigl(0,\,f_{\theta}(u_{A,i},x_{0,i})-f_{\theta}(u_{B,i},x_{0,i})+\delta\bigr),$ (6)
and minimize its empirical expectation over the dataset via stochastic gradient descent. This loss softly enforces the inequality $f_{\theta}(u_{A},x_{0})<f_{\theta}(u_{B},x_{0})-\delta$ whenever $A\succ B$. To handle the other case $B\succ A$, we post-process $\mathcal{D}$ so that the preferred component is always the first one in the tuple.
To guarantee computational tractability of the downstream optimization, the model $f_{\theta}$ is not learned as a generic neural network but is shaped to produce a convex quadratic objective of the form
@@
where the entries of $H(x_{0})\succeq 0$ and $q(x_{0})$ are the output of a standard deep neural network. In particular, the NN provides as output the set of parameters $\theta=\{\theta_{i}\}_{i=1}^{n^{2}+n}$, and the first $n^{2}$ parameters fill the matrix $H$, while the last $n$ parameters fill $q$ (please refer to (Abdufattokhov et al., 2021) for further details on the approach). This choice guarantees that the optimization problem (5), as $x_{0}$ is fixed, can be solved with standard quadratic (mixed-integer) solvers without altering the constraint structure or the solver stack and possibly preserving convexity.
To train $f_{\theta}(u,x_{0})$, we employ the well-known Siamese training scheme: two identical copies of the same network $f_{\theta}(\cdot,x_{0})$ are evaluated (Chicco, 2021), respectively, on the paired actions $(u_{A},x_{0})$ and $(u_{B},x_{0})$, and trained through the hinge ranking loss (6). This architecture is not a different model class but a numerical device that computes both sides of each preference pair through the same parameterized function, ensuring that learning directly enforces the desired pairwise ordering.
@@
Although input–convex neural architectures (ICNNs) could in principle learn more general convex surrogates objectives, optimizing over these functions would make (5) fully non-linear. While this may be manageable in a continuous setting, it becomes highly problematic in mixed-integer formulations, where efficient MINLP solvers are rare and often unreliable. Our parametrization is therefore intentionally designed to avoid this issue.
@@ 3.1 Reducing Training Complexity via PCA Compression
The overall workflow of the demonstrator is summarized in Fig. 1.
Feasible Solutions GenerationUser Preference AcquisitionPairwise Dataset ConstructionSiamese DNN Quadratic Cost GeneratorLearned Convex Quadratic Objective $\tfrac{1}{2}u^{\top}H(x_{0})u+q(x_{0})^{\top}u$Reintegration into Solver
Figure 1: Schematic overview of the demonstrator workflow. Feasible suboptimal solutions are generated via biased optimization, scored by a simulated user function, and used to train a Siamese neural network that learns a convex quadratic objective consistent with user preferences.
It comprises five main stages: (i) generation of diverse feasible decisions via biased optimization; (ii) scoring of these solutions through a simulated user preference model; (iii) construction of a pairwise preference dataset; (iv) learning of a surrogate convex quadratic objective; and (v) reintegration of the learned objective into the original solver for fairness-aware optimization.
@@ 4.6 Experimental Evaluation
5 Conclusion
This paper proposed a data-driven framework to embed fairness in optimization-based Digital Twins through preference learning. We introduced a workflow that infers human-aligned surrogate objectives from pairwise preferences and realized it via a Siamese neural network that generates convex quadratic costs, ensuring tractability and solver compatibility. A proof-of-concept study on COVID-19 hospital resource allocation demonstrated that the learned objectives can improve perceived fairness without compromising nominal efficiency. Future work will extend this framework to scenarios involving multiple users with potentially diverging preferences, aiming to reconcile conflicting feedback signals and model collective fairness equilibria. Additional research will explore real human-in-the-loop and online learning scenarios.